### v3

# AI for Scientific Discovery is a Social Problem

Abstract
Artificial intelligence promises to accelerate scientific discovery, yet its benefits remain unevenly distributed. While technical obstacles such as scarce data, fragmented standards, and unequal access to computation are significant, we argue that the primary barriers are social and institutional. Narratives that defer progress to speculative “AI scientists,” the undervaluing of data and infrastructure contributions, misaligned incentives, and gaps between domain experts and machine learning researchers all constrain impact. We highlight four interconnected challenges: community dysfunction, research priorities misaligned with upstream needs, data fragmentation, and infrastructure inequities. We argue that their roots lie in cultural and organizational practices. Addressing them requires not only technical innovation but also intentional community-building, cross-disciplinary education, shared benchmarks, and accessible infrastructure. We call for reframing AI for science as a collective social project, where sustainable collaboration and equitable participation are treated as prerequisites for technical progress.
1 Introduction
Artificial intelligence promises to accelerate scientific discovery, yet its benefits remain concentrated rather than democratized. While breakthrough applications like AlphaFold [49] and the Materials Project [44] demonstrate AI’s transformative potential, current development remains largely inaccessible to most researchers, requiring computational resources, specialized datasets, and interdisciplinary expertise that few institutions can match [72]. This concentration limits research diversity and undermines the collaborative foundations essential for robust scientific progress [26].
The challenge extends beyond resource scarcity to a fundamental mismatch between AI paradigms and scientific needs. Most AI development optimizes for predictive accuracy on large, homogeneous datasets, while science demands understanding from high-dimensional, low-sample-size data with complex relationships [10]. When AI tools prioritize automation over augmentation, they risk short-circuiting the human understanding essential for scientific progress [54], creating a production-progress paradox where increased output does not translate to deeper insight [50].
This paper examines four critical barriers preventing AI democratization in science: community dysfunction that undermines collaboration, misaligned research priorities targeting narrow applications over upstream computational bottlenecks, data fragmentation due to incompatible standards [38], and infrastructure inequities concentrating power within privileged institutions [4]. While these manifest as technical challenges, their root causes are fundamentally social and institutional: including undervaluing data contributions, educational gaps preventing cross-disciplinary engagement, and absent community consensus on shared priorities. We then propose corresponding solutions: strengthening cross-disciplinary collaboration and education, structuring upstream challenges through shared benchmarks, standardizing scientific data practices, and building sustainable community-owned infrastructure. Our analysis demonstrates that democratizing AI for science requires treating it as a collective social project where equitable participation is a prerequisite for technical progress.
2 Barriers to Scientific Achievement
2.1 Barrier One: Community Dysfunction
The dysfunction within the AI for science community manifests through harmful narratives, misaligned incentives, and communication breakdowns that prevent effective collaboration. These reflect deeper cultural issues about how the community values different types of contributions, frames the purpose of AI in science, and structures interactions between disciplines.
The AI Scientist Myth
Figure 1: Figure from  [85] contrasts the true Newtonian forces (left) and the predicted forces (right) learned by a transformer-based foundation model with high accuracy in predicting planetary trajectories. Though it performs well on the task it was fine-tuned for, it has not learned an inductive bias toward true Newtonian mechanics. General purpose models do not necessarily aid in specific scientific understanding.
It has become popular to suggest that artificial general intelligence (AGI) is imminent and, that once it arrives, will solve the vast majority of our scientific problems [42]. Many efforts in the machine learning for science community have also focused on the development of so-called “AI scientists”, which have had results ranging from disappointing to actively harmful [27]. While these systems may eventually provide valuable support as co-pilot tools, particularly in retrieving accurate information and assisting with routine tasks, they are often framed as a future replacement for the human scientist.
This narrative is ultimately counterproductive to genuine scientific progress [15]. First, there is little evidence that AGI is imminent, let alone consensus on what the term means, and waiting for it risks delaying advances that could be achieved with current methods. Second, it devalues the expertise and contributions of human scientists, whose knowledge and creativity remain essential to discovery [45]. Third, it oversimplifies the inherent complexity of scientific practice, which depends not only on predictive accuracy but also on careful verification, contextualization, and theoretical integration [94]. Recent work by Luo et al.  [64] demonstrates that “AI scientists” are prone to failure modes, such as inappropriate benchmark selection, data leakage, metric misuse, and post-hoc selection bias, which are likely to be overlooked when evaluating only the final output. Finally, the vision of the fully autonomous “AI scientist” obscures the central purpose of science, which is not only to produce solutions but also to cultivate human understanding [24].
Rewarding Long-Lasting Impact
Another counterproductive tendency within the current research ecosystem is the undervaluing of contributions to data and infrastructure in hiring, publicity, and tenure evaluations [34]. High-quality datasets, particularly those with well-curated metadata, often have far greater long-term impact than individual model contributions. Most models are rapidly superseded by marginally improved variants, and their influence diminishes within a short time frame. By contrast, datasets and infrastructure frequently underpin entire lines of research and remain useful for decades [73]. The effective half-life of data is therefore likely to be orders of magnitude longer than that of models. Recognizing and rewarding these contributions is essential for building a sustainable foundation for scientific progress with machine learning.
Technical Communication Gaps
Collaboration failures often stem from fundamental differences in how fields approach problem formulation and validation [55]. Domain scientists prioritize mechanistic understanding and experimental validation, while ML researchers focus on predictive performance and computational efficiency. Depending on the field, they may also rely on different tools and software practice, which further reinforces disciplinary boundaries. These different objectives lead to miscommunication about success criteria, acceptable trade-offs, and validation approaches.
The terminology gap exacerbates these issues. Domain scientists may describe problems using field-specific concepts that don’t map directly to standard ML problem formulations. ML researchers may propose solutions that seem promising from a computational perspective but violate fundamental domain constraints or assumptions. For example, a machine learning model might achieve excellent accuracy in predicting rainfall patterns but do so by overfitting to seasonal cycles in the training data. Without incorporating physical constraints from climate dynamics, these predictions fail under shifting climate regimes and provide little scientific insight.
Insufficient Educational Materials
A critical challenge lies in the lack of educational resources tailored for interdisciplinary research. Machine learning practitioners seeking to apply methods in biology, chemistry, and physics lack primers designed for technically mature but domain-naive audiences, while domain scientists frequently lack training in computational and ML methods. This bidirectional gap reflects the fact that most researchers trained in traditional sciences without computing requirements, often leading them to reinvent existing solutions [80]. Just as mathematics and statistics are fundamental to scientific training, machine learning should now be treated as essential for domain scientists. Without this foundation, collaborative efforts suffer as each group struggles to recognize what the other can contribute [13].
2.2 Barrier Two: Misaligned Research Priorities
The scientific community’s focus on narrow, domain specific applications over upstream computational challenges reveals how academic reward systems shape research agendas. Publication pressure, grant cycles, and disciplinary boundaries create powerful incentives that fragment effort across hundreds of domain-specific problems rather than mobilizing collective action around high-impact computational bottlenecks.
Identifying Upstream Computational Bottlenecks
Certain computational challenges appear across numerous scientific domains, representing high-leverage targets for AI development. Fast density functional theory (DFT) calculations underpin materials discovery, catalysis research, and solid-state physics [35]. Efficient partial differential equation (PDE) solvers enable real-time fluid dynamics, heat transfer modeling, and climate simulations [52]. Long-timescale molecular dynamics captures biochemical processes and materials aging across chemistry, biology, and engineering.
The Critical Assessment of protein Structure Prediction (CASP) illustrates how structuring an upstream challenge can catalyze broad progress. By framing protein structure prediction as a recurring community benchmark with standardized evaluation metrics, CASP identified the key computational bottleneck and created a venue for sustained methodological innovation [68]. This groundwork ultimately enabled breakthroughs such as AlphaFold, but the critical contribution was the clear problem formulation and consistent evaluation regime that mobilized an entire research community [49]. By focusing collective effort on a shared upstream challenge, CASP multiplied downstream impact compared to piecemeal advances on narrow protein engineering tasks.
Examples of High-Impact Computational Problems
Several computational challenges represent similar opportunities for broad impact, for example:
  • Efficient PDE solvers: Generalizable solvers that work across different equations and report confidence when extrapolating beyond training data [86].
  • Multi-scale coupling: Linking atomistic simulations to continuum behavior for complex systems spanning multiple physical scales [25].
  • High-dimensional sampling: Underpinning uncertainty quantification and Bayesian inference across experimental sciences [14].
  • Quantum many-body methods: Tackling strongly correlated systems in condensed matter physics and quantum chemistry [79].
  • Fast materials property prediction: Enabling high-throughput screening for battery materials, catalysts, and semiconductors [90].
The challenge lies in identifying these upstream problems within specific research communities. Domain scientists often focus on their specialized applications without recognizing broader computational patterns. Machine learning researchers may lack the domain knowledge to identify which computational bottlenecks have the widest applicability.
Figure 2: Scientific data tokenization challenges. Current AI architectures excel at predicting text tokens (top) but struggle with complex scientific datasets like multi-omics data (bottom), which lack clear tokenization strategies and exhibit low predictive capacity despite their richness and scale.
2.3 Barrier Three: Data Challenges
Data fragmentation in science refers to the scattering of research data across disconnected sources, stored in incompatible formats without shared standards or effective indexing, which makes discovery, integration, and reuse difficult. It stems from a community that has failed to prioritize interoperability and sharing despite generating vast quantities of valuable datasets. Researchers hoard data in proprietary formats, funding agencies don’t require standardization, and institutions provide no career rewards for the tedious work of data curation and harmonization. The result is a landscape where most scientific datasets remain locked in incompatible silos, while the high-dimensional, low-sample-size nature of scientific data makes these coordination failures even more costly for AI development.
The High-Dimensional, Low-Sample-Size Problem
Consider protein structure prediction: while AlphaFold succeeded with millions of protein sequences, most scientific domains lack equivalent data abundance [49]. A typical cancer genomics study might have 10,000 features (genes) but only 100 patients. Whole slide imaging (WSI) contains billions of pixels with complex spatial hierarchies, yet training sets rarely exceed thousands of samples [63]. Current transformer architectures, designed for text where small context windows suffice, generally fail to capture the extended spatial relationships essential for scientific understanding [19]. As illustrated in Figure 2, while AI models excel at predicting text through established tokenization strategies, scientific datasets like multi-omics datasets, including transcriptomics, proteomics, and metabolomics, contain complex spatial, temporal, and multimodal relationships that resist straightforward tokenization approaches [6], resulting in poor predictive performance despite the large scale and richness of the underlying data.
The context window limitation becomes stark in biological systems. A transformer trained on protein sequences might capture local amino acid patterns [28], but understanding protein function requires considering interactions across the entire three-dimensional structure—often requiring 1000x larger context windows than current models support. Random autoregressive rollout works for text generation because local context provides meaningful signal, but protein folding depends on long-range interactions that sequence models are inherently limited in their ability to represent, and require specific architecture modifications [82].
Data Fragmentation and Standardization Challenges
Beyond architectural limitations, scientific data suffers from severe fragmentation that prevents collaborative model development. Research organizations typically manage over 100 distinct data sources, with 30% handling more than 1,000 sources [75]. For instance, a 2020 survey found that data scientists spend about 45% of their time on data preparation tasks, including loading and cleaning data [93]. Domain-specific formats, inconsistent metadata standards, and incompatible experimental protocols create artificial barriers that fragment knowledge and limit collaborative development.
The cancer research field exemplifies these challenges. Harmonizing just nine Cancer Imaging Archive files required 329.5 hours over six months, identifying only 41 overlapping fields across three or more files [9]. This labor-intensive process must be repeated for each new research initiative, creating insurmountable barriers for smaller institutions lacking dedicated data curation resources.
Similar problems arise outside of the life sciences. In astroparticle physics, experiments like KASCADE and TAIGA produced distributed datasets in heterogeneous, experiment-specific formats, such as ROOT trees, custom binaries, and ASCII logs, that required specialized software just to parse [43]. Because files were large and monolithic, extracting subsets (e.g., high-energy showers) often meant loading gigabytes of irrelevant data. Data were also scattered across institutions with limited metadata, creating significant computational and logistical overhead before scientific analysis could even begin [56].
Data Format and Evaluation Incompatibilities
Domain scientists typically generate datasets optimized for human interpretation rather than machine learning training. Experimental data often comes in specialized formats with implicit assumptions about preprocessing, normalization, and quality control that ML researchers must reverse-engineer. Different fields use incompatible evaluation metrics, making it difficult to assess whether ML approaches actually improve upon domain-specific baselines.
For example, computational biology datasets often require complex preprocessing pipelines to handle batch effects, missing values, and experimental artifacts [65]. Without domain expertise, ML researchers may train models on processed data without understanding the underlying experimental limitations, leading to overfitting to technical artifacts rather than biological signal.
2.4 Barrier Four: Infrastructure Inequity
Infrastructure inequity in scientific computing reflects broader patterns of resource concentration and institutional privilege that go beyond technical necessity. The social dynamics of resource allocation: who gets priority, how decisions are made, and which institutions are considered “deserving”, ultimately determine whether AI tools democratize or further centralize scientific capability. The challenge lies in building systems that work effectively across different resource constraints while maintaining scientific quality and enabling collaborative development.
Computational Resource Disparities
Academic researchers face substantial barriers accessing computational resources adequate for competitive AI development. Nature’s survey found that 66% of scientists rated their computing satisfaction as three out of five or less, with only 10% having access to state-of-the-art hardware [57]. Wait times commonly extend days to weeks, while complex application processes favor institutions with existing technical infrastructure.
Global disparities are even more pronounced. Only 5% of Africa’s AI research community has access to computational power needed for complex AI tasks, while 95% rely on limited free tools [84]. African researchers face six-day iteration cycles compared to 30-minute cycles available in G7 countries, creating fundamental inequities in research capability that no amount of algorithmic innovation can overcome.
Computational Infrastructure Misalignments
Many researchers in computational domain sciences are also members of the high-performance computing (HPC) community, which has long operated under a different set of constraints and practices. HPC systems are in general optimized for large sequential reads and writes (e.g., simulation checkpoints), not the millions of tiny files typical in modern ML datasets [60]. For instance, system administrators may be displeased to see training directly on ImageNet-style datasets where every image is a small, individual file, since this can place enormous strain on the file system’s metadata servers by forcing constant indexing. This divergence in infrastructure expectations can create additional friction: ML researchers may be surprised when practices that are routine in their community are discouraged in some HPC settings, which can complicate collaboration on shared compute resources or with shared tooling.
3 Solutions and Implementation Strategies
If these barriers were addressed, the scientific community could achieve outcomes that remain out of reach today. Researchers across domains could build foundation models that integrate knowledge from biology, chemistry, physics and materials science rather than maintaining fragmented efforts. Experimental laboratories of all sizes could run rapid experiment–model loops that reduce iteration cycles from weeks to hours and accelerate discovery across disciplines. Scientists could rely on trustworthy AI co-pilots that highlight inconsistencies, suggest experiments and contextualize results. These systems would be grounded in robust data and reproducible infrastructure. Communities could rally around shared grand challenge benchmarks to drive progress on upstream computational problems. Most importantly, these advances would become accessible to a global community of researchers and would enable contributions from institutions and regions that are currently excluded by resource and infrastructure disparities.
3.1 Solution One: Strengthening Collaboration and Education Across Communities
Standardized Interfaces and APIs
Technical collaboration barriers can be addressed through standardized interfaces that abstract implementation details while preserving domain-specific requirements. Machine learning libraries should provide scientific computing interfaces that handle common preprocessing steps, domain-specific evaluation metrics, and uncertainty quantification automatically [92, 71, 74].
Successful examples include scikit-image for microscopy [87] and astropy for astronomical data [21], which demonstrate how standardized interfaces can expose simple methods to domain scientists while handling complex domain requirements like irregular sampling, missing data, and measurement uncertainties. Version control and reproducibility tools adapted for scientific workflows can facilitate collaboration by tracking code and data provenance [48], though tools like DVC and MLflow need extension to handle domain-specific requirements like experimental metadata and regulatory compliance [78].
Collaborative Development Platforms and Training
Successful technical collaboration requires platforms that enable iterative development between domain experts and ML researchers. While platforms like Hugging Face demonstrate effective community-driven development, scientific applications require additional features including uncertainty quantification, domain-specific evaluation metrics, and validation against established scientific benchmarks [17]. Building scientific AI repositories that incorporate these requirements by default would substantially lower barriers to collaboration [61].
This requires training programs that build hybrid expertise rather than attempting to turn domain scientists into ML specialists or vice versa [3]. The community should develop specialized roles for scientific AI practitioners who can bridge both domains effectively [51]. Online educational resources should prioritize practical skills for scientific AI applications—preparing domain datasets, validating ML models against scientific knowledge, and interpreting results in context [62]. These resources should emphasize understanding over automation, enabling researchers to use AI tools responsibly rather than as opaque black boxes.
3.2 Solution Two: Structuring Upstream Challenges and Shared Benchmarks
Community Mechanisms for Problem Identification
The research community needs systematic approaches to identify computational bottlenecks with broad applicability through dedicated conference sessions and cross-disciplinary workshops. Prize competitions demonstrate effective models for community-driven problem identification: the Vesuvius Challenge [2], DREAM Challenges [47] have engaged over 30,000 participants across sixty competitions producing 105+ publications, while DrivenData competitions [29] have provided \$4.8M+ in prizes across 245,000 submissions advancing climate, health, and education solutions. These competitions attract diverse expertise to problems individual research groups cannot tackle alone, using hybrid competition-collaboration models that prevent knowledge hoarding while maintaining competitive motivation.
Research funding programs should explicitly prioritize computational challenges with broad applicability rather than funding numerous domain-specific applications, as concentrated effort on upstream problems can achieve greater total impact. The community should develop frameworks for assessing the potential downstream impact of computational advances to guide strategic resource allocation.
Building Reusable Computational Infrastructure
Successful upstream solutions require careful attention to generalizability and reusability through open-source development practices, well-documented APIs, comprehensive test suites, and modular architectures that enable researchers to build upon and extend computational tools [30]. The community should prioritize sustainability through proper software engineering practices rather than treating research code as disposable prototypes, while ensuring that reproducible science is valued and rewarded in hiring, tenure evaluations, and citation counts.
Benchmark datasets and evaluation metrics for upstream problems should focus on computational efficiency, accuracy across diverse problem instances, and uncertainty quantification rather than performance on narrow tasks [91]. For the scientific workforce to coalesce around such challenges, domain leaders must take responsibility for defining them and actively engaging their communities via easy to use open benchmark creation methods [81]. The most effective step individual researchers can take is to help articulate and participate in community benchmarks that frame upstream computational challenges as shared problems rather than siloed pursuits.
3.3 Solution Three: Standardizing and Curating Scientific Data for Broad Reuse
Community-Driven Data Standardization
Successful democratization of scientific data requires practical, widely adopted standards that balance usability with scientific integrity. Simple, universally compatible formats like CSV or Parquet often lower barriers to collaboration more effectively than complex, domain-specific schemas. Early bioinformatics efforts, such as MAGE-ML [83], illustrate the pitfalls of overly complex standards: despite technical flexibility, its complexity hindered interoperability, leading the community to adopt the simpler spreadsheet-based MAGE-TAB format [77, 76]. Similarly, attempts to harmonize nine Cancer Imaging Archive files required 329.5 hours of manual effort that could have been largely avoided with standardized schemas [20].
The Protein Data Bank (PDB) exemplifies large-scale, community-driven data democratization. Starting with seven structures in 1971, it now houses over 230,000 three-dimensional protein structures with 10 million daily downloads [8, 12]. Its success stems from standardized formats (PDB/mmCIF), expert curation, open access principles, and sustained funding. Similar approaches—minimal metadata standards, conversion tools for legacy formats, and validation pipelines—can enable other domains to achieve scalable reuse and collaborative development [73, 92, 39].
Even modest adoption of open, accessible formats can have outsized impact. Open-source tools like Pandas, Polars, and Apache Arrow provide robust cross-platform support for scientific data processing, facilitating over a billion computations daily [88]. The growth of repositories such as Zenodo, which hosts over 2 million datasets following simple Dublin Core metadata standards, underscores how accessible standards accelerate data sharing and democratization [31].
Developing Architectures for Scientific Data
The high-dimensional, low-sample-size problem requires architectural approaches that extend beyond current transformer models. Progress has been made in developing architectures tailored for scientific data. Graph neural networks show promise for molecular and materials applications by explicitly modeling spatial relationships, achieving 20% improvements in catalyst activity prediction over traditional descriptors [37, 36, 18]. Hierarchical attention mechanisms have been proposed to address the long-range interaction problem in biological systems, enabling analysis of gigapixel pathology images with spatial hierarchies spanning six orders of magnitude [16, 19, 41]. Physics-informed neural networks incorporate domain knowledge directly into model design, reducing simulation errors by 2-3 orders of magnitude in fluid dynamics applications while requiring 100x less training data [59, 89, 53]. State space models like Mamba [40] offer promising alternatives for handling long sequences common in scientific time series, achieving linear scaling with sequence length compared to quadratic complexity in transformers. Yet better inductive biases are still needed, and embedding them effectively into architectures remains non-trivial [85, 58]. Deep knowledge of both machine learning and the relevant domain is essential for success. Further, each of the research areas mentioned above lacks robust benchmarking and consequently has struggled to gain sustained traction [11].
However, architectural innovation alone is insufficient without addressing the sample efficiency problem. Few-shot learning approaches, meta-learning techniques, and transfer learning from simulation data offer promising directions, with recent demonstrations showing 90% accuracy achievable using only 50 experimental samples when pre-trained on 10 million simulated materials [67, 7, 70, 89, 5]. The community should prioritize developing foundation models that capture general scientific principles and can be fine-tuned for specific applications with limited data, requiring coordinated dataset construction efforts spanning multiple institutions and decades of sustained funding commitment.
3.4 Solution Four: Building Accessible and Sustainable Infrastructure
Efficient Model Sharing and Deployment
Democratization requires infrastructure for sharing not just trained models but entire scientific AI pipelines. Scientific models often require complex preprocessing, specialized evaluation, and careful uncertainty quantification that generic model repositories don’t support adequately.
Model efficiency becomes critical for equitable access. Quantization, pruning, and knowledge distillation can dramatically reduce computational requirements while preserving scientific accuracy [23], but these techniques need adaptation for scientific applications rather than simply borrowing consumer-focused optimizations [66]. For instance, scientific models must maintain calibrated uncertainty estimates even after compression. Edge deployment strategies are particularly important for field research, remote laboratories, and institutions with limited computational infrastructure [33].
Community-Owned Infrastructure and Sustainable Funding
Sustainable infrastructure depends on community ownership rather than reliance on commercial platforms or individual institutions. Open source projects like Jupyter, Conda, and the Scientific Python ecosystem illustrate successful community-driven models that serve broad research needs. Scientific AI systems, however, require additional components such as provenance tracking for reproducibility, specialized metadata schemas, and support for long-running computations.
Building this infrastructure is critical, but so is investment in outreach and dissemination. Valuable resources already exist—such as free compute, storage, and API credits from NAIRR [69], OSPool [1], and Hugging Face [32]—yet often go unused because potential users remain unaware of them. Without active education and visibility, infrastructure risks being underutilized.
The central challenge is sustainable funding that supports both long-term development and the communication strategies needed for adoption. Infrastructure takes years of effort, but traditional research funding favors novelty over maintenance [22]. Addressing this gap will require mechanisms that combine support for infrastructure and dissemination, through consortium models, government investment, or hybrid public–private partnerships [46].
4 Discussion
Our analysis underscores a central challenge for AI in science: widespread adoption alone does not guarantee deeper understanding or more equitable progress. Many barriers are social rather than technical, including undervaluing contributions to data and infrastructure, fragmented interdisciplinary collaboration, and incentive structures that favor narrow advances over broad impact.
Addressing these challenges requires integrating technical improvements with community-centered strategies. Strengthening collaboration and education, structuring upstream benchmarks, standardizing data practices, and building accessible infrastructure can create a foundation for shared, sustainable progress. Realizing AI’s transformative potential depends as much on aligning community priorities and incentives as on advancing algorithms and hardware.
Limitations
The proposed solutions require coordinated funding, long-term institutional commitment, and policy support, particularly for infrastructure initiatives, which benefit from stability beyond the lifespan of individual grants. Achieving widespread adoption also depends on overcoming community inertia and entrenched tool preferences. Success requires technical excellence combined with sustained engagement, trust-building, and clear migration pathways. While the primary audience for this paper is the scientific community, implementing a truly holistic solution also requires coordinated action from policymakers, AI researchers, funders, and other stakeholders.
Future Work
International collaboration can reduce disparities in computational access and support equitable participation in scientific AI, but must address data sovereignty, technology transfer, and fair governance.
Future work should establish metrics for democratization that go beyond adoption counts to include contributing institutions, geographic distribution, and capacity gains at under-resourced organizations. Such measures will help ensure progress is genuine, widely shared, and aligned with AI’s transformative potential.